OpenStudy (anonymous):

Given the function k described by k(x)= x+4, find the following: k(0)=?

OpenStudy (anonymous):

Just let x=0. So... k(0)=(0)+4 k(0)=4 And there is your answer :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

k(-1) =?

OpenStudy (anonymous):

No problem :) For k=(-1), its the same idea... k(-1)=(-1)+4 k(-1)=-1+4 k(-1)=3

OpenStudy (anonymous):

k(-13) =?

OpenStudy (anonymous):

k(-13)=(-13)+4 k(-13)=-9

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

uh oh...k(12)=?

OpenStudy (anonymous):

k(12)=(12)+4 k(12)=16

OpenStudy (anonymous):

Whoo Hoo!!!

OpenStudy (anonymous):

One more.... k(c+6) = ?

OpenStudy (anonymous):

k(c+6)=(c+6)+4 k(c+6)=c+10

OpenStudy (anonymous):

In this case you don't get a numerical value, since you are replacing the variable x with another variable.

OpenStudy (anonymous):

You were right every time. Damn you are smart!

OpenStudy (anonymous):

It's simple substitution :) You'll probably get the hang of it quickly.